1. Calculate the price of Zachary Corp.'s common stock with the Dividend Growth Model given that dividends were $2.30 per share last year, are expected to grow 3% per year for the foreseeable future, and investors' required return is 14%. Calculate to 4 decimal places.
  2. A.$21.54B.$20.91C.$12.48D.$14.35
1
Expert's answer
2021-03-03T10:22:23-0500

Current stock price:


"P=\frac{D}{r-g}"

"g" - growth rate expected for dividends

"r" - rate of return

"D" - value of next year's dividends


"P=\frac{2.3}{0.14-0.03}=20.91" (Answer B)